OpenStudy (anonymous):

find exact value of cos(2theta) given that sintheta= -8/17 and theta is in Quad III

OpenStudy (anonymous):

\[\huge \cos(2\theta) = 1- 2 \sin^2 \theta =1-2 (\frac{-8}{17})^2\] \[\huge =1-2 (\frac{64}{289})=1-\frac{128}{289} = \frac{289-128}{289}\] \[\huge \cos (2 \theta) = \frac{161}{289}\]

OpenStudy (anonymous):

thanks so much!